Practice Brief Exercise 15--05

Bramble Corporation has outstanding 47,000 shares of $50 par value, 5% preferred stock and 336,000 shares of $1 par value common stock. The preferred stock was issued in July 2017, and no dividends were declared in 2017 or 2018. In 2019, Bramble declares a cash dividend of $482,000.

How will the dividend be shared by common and preferred stockholders if the preferred is noncumulative?
Preferred Stockholders $enter a dollar amount

Common Stockholders $enter a dollar amount

How will the dividend be shared by common and preferred stockholders if the preferred is cumulative?
Preferred Stockholders $enter a dollar amount

Common Stockholders $enter a dollar amount
Answer #1

Preferred dividend= 47000*$50*5%= $117500

1) How will the dividend be shared by common and preferred stockholders if the preferred is noncumulative?

Preferred dividend= $117500

Common dividend= Total dividend-Preferred dividend

= $482000-117500= $364500

2) How will the dividend be shared by common and preferred stockholders if the preferred is cumulative?

In case of cumulative preferred shares, the shareholders will get the dividend of the previous years even if the company does not declare dividend. As the company did not declare dividend in 2017 and 2018, the preferred shareholders will get the dividend of 2017 and 2018 along with 2019.

Preferred dividend= Dividend for 2017+Dividend for 2018+Dividend for 2019

= $117500+117500+117500= $352500

Common dividend= Total dividend-Preferred dividend

= $482000-352500= $129500